Woidtke Manufacturing's stock currently sells for $38 a share. The stock just paid a dividend of $1.75 a share (i.e., D0 = $1.75), and the dividend is expected to grow forever at a constant rate of 10% a year. What stock price is expected 1 year from now? Round your answer to the nearest cent.
$  

What is the estimated required rate of return on Woidtke's stock? Do not round intermediate calculations. Round the answer to three decimal places.
     %
Answer #1

P1 = P0(1 + g) = $38(1 + 0.10) = $41.80

r = D1/P0 + g = [$1.75(1.10) / $38] + 0.10 = 0.1507 or 15.07%